Trente Corp has identified two activities in its activity based costing system: setups and production. The activity overhead rates for these two activities are $321 per setup and $1.31 per direct labor hour, respectively. To make product X5, Trente incurred $623 in direct materials costs, $124 in direct labor costs, and used 3 setups and 29 direct labor hours. The total overhead assigned to product X5 would be:
Answer #1

Total overhead Assigned to Product X5

Setup Cost = 3*321 = $ 963

Production Cost = 1.31 *29 = $ 37.99

Overhead Assigned = 963+37.99 = $1000.99 (or) $1001 (Answer)